• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name markowitz

    Log entry time 23:57:05 on June23,2000

    Entry number 44741

    keyword=swing shift summary



    Started shift with accelerator tuning going on.
    Experiment doing production running at thcm = -50

    16:00 Injector tuning underway. I>60uA was tripping off the fsd
    17:20 Mcc will send tuning beam into the hall.
    18:00 setting up for moller run (by Eugene and Victor)
    19:00 still waiting for beam for the moller running.
    19:29 Moller done. Mcc says dipole before running moller was off and ask
    if we want it off now. Eugene says it should be ok to turn it back off.
    19:40 Mcc reports RF problems. Beam should be back soon.
    20:30 mcc says a cavity won't stay locked and will need bypassed beam to be gone
    for about 10 minutes.
    20:45 mcc has bypassed the rogue cavity (which is now downgraded to a
    "cavity-of-concern") and is ready to deliver beam again
    23:30 rf trip. mcc will do a qe scan when they come back.
    23:35 beam back

    run events start stop comments
    2477 0.8M 19:55 20:27 beam unstable for early part of run
    2478 1.3M 20:45 21:25 still unstable rf beam
    2479 1.3M 21:26 22:05 beam stable
    2480 1.3 M 22:05 22:48 beam rf unstable
    2481 1.3M 22:49 23:27 beam more stable
    2482 underway

    through run 2481 we have 5.35 C of charge.